# Boosting Continuous Control with Consistency Policy

### v1

@@ Abstract.
Due to its training stability and strong expression, the diffusion model has attracted considerable attention in offline reinforcement learning. However, several challenges have also come with it: 1) The demand for a large number of diffusion steps makes the diffusion-model-based methods time inefficient and limits their applications in real-time control; 2) How to achieve policy improvement with accurate guidance for diffusion model-based policy is still an open problem. Inspired by the consistency model, we propose a novel time-efficiency method named Consistency Policy with Q-Learning (CPQL), which derives action from noise by a single step. By establishing a mapping from the reverse diffusion trajectories to the desired policy, we simultaneously address the issues of time efficiency and inaccurate guidance when updating diffusion model-based policy with the learned Q-function. We demonstrate that CPQL can achieve policy improvement with accurate guidance for offline reinforcement learning, and can be seamlessly extended for online RL tasks. Experimental results indicate that CPQL achieves new state-of-the-art performance on 11 offline and 21 online tasks, significantly improving inference speed by nearly 45 times compared to Diffusion-QL. We will release our code later.
@@ 3.2. Diffusion Policy
In this section, we describe how to use the diffusion model to build the expected policy of Eq.(2) and define this policy as diffusion policy. Before formally introducing this process, we first declare that there are two different types of timesteps in the following sections. Rather than using subscripts $t$ denoting the trajectory timesteps, we use superscripts $k\in[0,K]$ to denote the diffusion timesteps.
Diffusion model starts by diffusing $p_{data}(a)$ with a stochastic differential equation (SDE):
@@
When the diffusion model is used to represent the expected policy directly, the challenge lies in estimating the score function because it’s impossible to obtain data from the optimal policy straightforwardly, making estimating the score function intractable. Nevertheless, we can use the diffusion process through reference policy with guidance to estimate the score function Dhariwal and Nichol (2021). With Eq. (2), we can derive the following score function with guidance for the solution of Eq. (2):
@@ 4.1. Consistency Policy
where $a^{k}\sim\mathcal{N}(0,kI)$. The $F_{\theta}(a^{k},k|s_{t})$ is a trainable deep neural network that takes the state $s_{t}$ as an input condition and outputs an action of the same dimensionality as $a^{k}$. $c_{skip}$ and $c_{out}$ are differentiable functions such that $c_{skip}(\epsilon)=1$, and $c_{out}(\epsilon)=0$ to ensure the consistency policy is differentiable at $k=\epsilon$ if $F_{\theta}(a^{k},k|s)$ is differentiable, which is critical for training process described later. We stop solving the reverse process at $k=\epsilon$, where $\epsilon$ is a small positive constant to avoid numerical instability. The sampled action $\hat{a}^{\epsilon}\sim\pi_{\theta}(a|s)$ of the consistency policy is used for controlling the agent. The relationship between the forward diffusion process represented by the ODE trajectories and consistency policy is shown in Figure 1.
@@ 5. Experiments
In offline tasks, we evaluate the methods on two domains of the D4RL benchmark. We compare CPQL with current methods that achieve SOTA performance. These methods includes Q-learning with policy constraints such as TD3+BC Fujimoto and Gu (2021), implicit Q-learning such as IQL Kostrikov et al. (2022), and diffusion-model-based methods, such as Diffuser Janner et al. (2022), DecisionDiffuser (DD) Ajay et al. (2023), SfBC Chen et al. (2023), IDQL Hansen-Estruch et al. (2023), Diffusion-QLWang et al. (2023b), EDP Kang et al. (2023), QGDO Lu et al. (2023b). Notably, Diffuser and DD employ the diffusion model to capture trajectory dynamics, while the remaining methods focus on modeling policy distributions.
@@ 5.1. Overall Results
In this section, we illustrate the overall evaluating results on the offline tasks (D4RL) and the online tasks (dm_control, Gym Mujoco). The results show that CPQL achieves very competitive performance on both offline and online tasks compared with current SOTA methods. Furthermore, we provide all training curves for CPQL and CPIQL in Appendix E.
@@ 5.1.1. Results on D4RL (Offline Tasks)
We first evaluate the performance of the methods on offline D4RL tasks. The experimental results are shown in Table 1. From the experimental results, we can see that the methods based on the diffusion model have significant advantages in performance compared with the unimodal distribution policy method (TD3+BC and IQL). Compared with Diffusion-QL and EDP, QGPO achieves excellent performance by using an accurate guided value function with contrastive loss. Our proposed CPQL avoids inaccurate guidance problems by modeling ODE trajectory mapping, thus achieving competitive results by approximately 4% improvement (compared with Diffusion-QL). By comparing CPQL and CPIQL, we can find that consistency policy can effectively constrain the sampling of OOD actions without the help of implicit Q-learning. Additionally, CPIQL outperforms IQL by a significant 16.5% on locomotion tasks, indicating that the strong representation ability of consistency policy can effectively achieve policy improvement with accurate diffusion guidance. In conclusion, CPQL achieves better guidance with value function while meeting constraints, thus making their performance outstanding on offline tasks.
@@ 5.1.2. Results on dm_control (Online Tasks)
Next, we evaluate the methods on 15 dm_control tasks. Part of the training curves and all experimental results are shown in Figure 2 and Table 2, respectively. As shown in Figure 2, CPQL has shown an improvement in both training stability and sample efficiency with better performance when compared to other algorithms under a small number of interactions with the environment (500 K). Also, the experimental results show that CPQL performance outperforms previous SOTA methods, including DreamerV3 on 15 tasks, especially on ”Cheetah Run” and ”Finger Turn Easy” tasks. On average, there has been an improvement of approximately 8% as compared to DreamerV3. In addition, CPQL directly models the desired policy and is not plagued by inaccurate diffusion guidance, resulting in better performance on complex tasks, with great representation ability for the policy. It is worth mentioning that DMPO and D4PG have adopted the distributional value function Bellemare et al. (2017), which has proved to achieve better performance than the expected value function. While CPQL only uses the expected value function, it has shown significant performance leadership, which also means the importance of exploration in RL, and the consistency model has great potential in RL.

### v2

@@ Abstract.
Due to its training stability and strong expression, the diffusion model has attracted considerable attention in offline reinforcement learning. However, several challenges have also come with it: 1) The demand for a large number of diffusion steps makes the diffusion-model-based methods time inefficient and limits their applications in real-time control; 2) How to achieve policy improvement with accurate guidance for diffusion model-based policy is still an open problem. Inspired by the consistency model, we propose a novel time-efficiency method named Consistency Policy with Q-Learning (CPQL), which derives action from noise by a single step. By establishing a mapping from the reverse diffusion trajectories to the desired policy, we simultaneously address the issues of time efficiency and inaccurate guidance when updating diffusion model-based policy with the learned Q-function. We demonstrate that CPQL can achieve policy improvement with accurate guidance for offline reinforcement learning, and can be seamlessly extended for online RL tasks. Experimental results indicate that CPQL achieves new state-of-the-art performance on 11 offline and 21 online tasks, significantly improving inference speed by nearly 45 times compared to Diffusion-QL. Code is available at https://github.com/cccedric/cpql.
@@ 3.2. Diffusion Policy
In this section, we define the expected policy of Eq.(2) as diffusion policy and describe how to use the diffusion model to build it. Before formally introducing this process, we first declare that there are two different types of timesteps in the following sections. Rather than using subscripts $t$ denoting the trajectory timesteps, we use superscripts $k\in[0,K]$ to denote the diffusion timesteps.
The diffusion model starts by diffusing $p_{data}(a)$ with a stochastic differential equation (SDE):
@@
When the diffusion model is used to represent the expected policy directly, the challenge lies in estimating the score function because it’s impossible to obtain data from the optimal policy straightforwardly, making the estimation intractable. Nevertheless, we can use the diffusion process through reference policy with guidance to estimate the score function Dhariwal and Nichol (2021). With Eq. (2), we can derive the following score function with guidance for the solution of Eq. (2):
@@ 4.1. Consistency Policy
where $a^{k}\sim\mathcal{N}(0,kI)$. The $F_{\theta}(a^{k},k|s)$ is a trainable network that takes the state $s_{t}$ as an condition and outputs an action of the same dimensionality as the input $a^{k}$. $c_{skip}$ and $c_{out}$ are differentiable functions such that $c_{skip}(\epsilon)=1$, and $c_{out}(\epsilon)=0$ to ensure the consistency policy is differentiable at $k=\epsilon$ if $F_{\theta}(a^{k},k|s)$ is differentiable, which is critical for training process described later. We stop solving the reverse process at $k=\epsilon$, where $\epsilon$ is a small positive constant to avoid numerical instability. The sampled action $\hat{a}^{\epsilon}\sim\pi_{\theta}(a|s)$ of the consistency policy is used for controlling the agent. The relationship between the forward diffusion process represented by the ODE trajectories and consistency policy is shown in Figure 1.
@@ 5. Experiments
For offline scenario, we evaluate on two domains of D4RL benchmark and compare CPQL with current methods that achieve SOTA performance, including Q-learning with policy constraints such as TD3+BC Fujimoto and Gu (2021), implicit Q-learning such as IQL Kostrikov et al. (2022), and diffusion-model-based methods, such as Diffuser Janner et al. (2022), DecisionDiffuser (DD) Ajay et al. (2023), SfBC Chen et al. (2023), IDQL Hansen-Estruch et al. (2023), Diffusion-QLWang et al. (2023b), EDP Kang et al. (2023), QGDO Lu et al. (2023b). Notably, Diffuser and DD employ the diffusion model to capture trajectory dynamics, while the remaining methods focus on modeling policy distributions.
@@ 5.1. Overall Results
In this section, we illustrate the overall results on the offline tasks (D4RL) and the online tasks (dm_control, Gym Mujoco), showing that CPQL achieves competitive performance on both offline and online tasks compared with current SOTA methods. All training curves for CPQL and CPIQL can be found in Appendix E.
@@ 5.1.1. Results on D4RL (Offline Tasks)
Firstly, we evaluate the performance of the methods on offline D4RL tasks. The experimental results are shown in Table 1. From the experimental results, we can see that the methods based on the diffusion model have significant advantages in performance compared with the unimodal distribution policy method (TD3+BC and IQL). Compared with Diffusion-QL and EDP, QGPO achieves excellent performance by using an accurate guided value function with contrastive loss. Our proposed CPQL avoids inaccurate guidance problems by modeling ODE trajectory mapping, thus achieving competitive results by approximately 4% improvement (compared with Diffusion-QL). By comparing CPQL and CPIQL, we can find that consistency policy can effectively constrain the sampling of OOD actions without the help of implicit Q-learning. Additionally, CPIQL outperforms IQL by a significant 16.5% on locomotion tasks, indicating that the strong representation ability of consistency policy can effectively achieve policy improvement with accurate diffusion guidance. In conclusion, CPQL achieves better guidance with value function while meeting constraints, thus making their performance outstanding on offline tasks.
@@ 5.1.2. Results on dm_control (Online Tasks)
Next, we evaluate the methods on 15 dm_control tasks. Part of the training curves and all experimental results are shown in Figure 2 and Table 2, respectively. As shown in Figure 2, CPQL has shown an improvement in both training stability and sample efficiency with better performance when compared to other algorithms under a small number of interactions with the environment (500 K). Also, the experimental results show that CPQL performance outperforms previous SOTA methods, including DreamerV3 on 15 tasks, especially on ”Cheetah Run” and ”Finger Turn Easy” tasks. On average, there has been an improvement of approximately 8% as compared to DreamerV3. In addition, CPQL directly models the policy and is not plagued by inaccurate diffusion guidance, resulting in better performance on complex tasks, with great representation ability for the policy. It is worth mentioning that DMPO and D4PG have adopted the distributional value function Bellemare et al. (2017), which has proved to achieve better performance than the expected value function. While CPQL only uses the expected value function, it has shown significant performance leadership, which also means the importance of exploration in RL, and the consistency model has great potential in RL.
@@ 6. Conclusion
{acks}
This work is supported by the National Natural Science Foundation of China (NSFC) under Grants No. 62136008, No. 62103409, the Strategic Priority Research Program of Chinese Academy of Sciences (CAS) under Grant XDA27030400 and in part by the International Partnership Program of the Chinese Academy of Sciences under Grant 104GJHZ2022013GC.